4y=x+32
5x+20y=0
Please use substitution method.

Found 2 solutions by checkley77, doukungfoo:
Answer by checkley77(12844)

4y=x+32 or x=4y-32
5x+20y=0
5(4y-32)+20y=0
20y-160+20y=0
40y=160
y=160/40
y=4 ans.
4*4=x+32
16=x+32
x=16-32
x=-16 ans.
Proof:
5*-16+20*4=0
-80+80=0
0=0

Answer by doukungfoo(195)

4y = x + 32
5x + 20y = 0
Ok to use substitution method you have to take one of the equations and set it equal to one of the variables. It doesn't matter which equation or which variable so I will use 4y=x+32 because to set it equal to x all you have to do is subtract 32 from both sides.
4y = x + 32
4y - 32 = x + 32 - 32
4y -32 = x
Now since we have shown that 4y - 32 = x
You can substitute 4y -32 for the x variable in the second equation, and solve for y.
5x + 20y = 0
5(4y-32) + 20y = 0 'replace x with 4y - 32
20y - 160 + 20y = 0 'multiply 5 times 4y and - 32
20y + 20y = 160 'add 160 to both sides
40y = 160 'combine like terms
y = 4 ' divide both sides by 40
Now that we have shown y to be equal to 4 we can substitute the y in the first equation with 4 and solve for x
4y = x + 32
4(4) = x + 32 ' replace y with 4
16 = x + 32 ' multiply 4 times 4
-16 = x ' subtract 32 from both sides
So for
4y = x + 32
5x + 20y = 0
x = -16 and y = 4
test solutions in both equations
4y = x + 32
4(4) = -16 + 32
16 = -16 + 32
16 = 16
AND
5x + 20y = 0
5(-16) + 20(4) = 0
-80 + 80 = 0
0 = 0